You are given the sample mean and the population standard deviation. Use this information to construct the​ 90% and​ 95% confidence intervals for the population mean. Interpret the results and compare the widths of the confidence intervals. If​ convenient, use technology to construct the confidence intervals. A random sample of 60 home theater systems has a mean price of ​$135.00 . Assume the population standard deviation is ​$15.90 .

A.) the 90% confidence interval is

B.) the 95% confidence interval is

C.) Which interval is wider?
Answer #1

Solution :

Given that,

= $135

= $15.90

n = 60

A)At 90% confidence level the z is ,

  = 1 - 90% = 1 - 0.90 = 0.1

/ 2 = 0.1 / 2 = 0.05

Z/2 = Z0.05 = 1.645

Margin of error = E = Z/2* ( /n)

= 1.645 * (15.90 / 60)

= 0.436

At 90% confidence interval estimate of the population mean is,

- E < < + E

135 - 0.436 < < 135 + 0.436

134.564 < < 135.436

(134.564,135.436 )

The 90% confidence interval is $134.564 to $135.436

B)At 95% confidence level the z is ,

  = 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

Margin of error = E = Z/2* ( /n)

= 1.96 * (15.90 / 60)

= 0.519

At 95% confidence interval estimate of the population mean is,

- E < < + E

135 - 0.519 < < 135 + 0.519

134.481 < < 135.519

(134.481,135.519 )

The 95% confidence interval is $134.481 to $135.519

C) The 95% interval is wider.